OpenStudy (anonymous):

What is the slope-intercept form of this equation? 8x + 6y = 12 A. Y = - 4/3x + 2 B. Y = -4/3x + 6 C. Y = 4/3x + 6 D. Y = 2/3x + 2

OpenStudy (theedwardsfamily):

WHat do you think @Nichole15 ? Do you know the formula?

OpenStudy (anonymous):

I think its A or B but I'm not sure

OpenStudy (theedwardsfamily):

ok. But do you know the formula for an equation like this

OpenStudy (theedwardsfamily):

*?

OpenStudy (anonymous):

No

OpenStudy (alexandervonhumboldt2):

8x + 6y = 12 first you need to subract 8x from both sides: 8x+6y-8x=12-8x 6y=12-8x divide both sides by 6: 6y/6=(12-8x)/6 y=-4/3x+2

OpenStudy (theedwardsfamily):

This is y=mx +b

OpenStudy (alexandervonhumboldt2):

yeah it is

OpenStudy (theedwardsfamily):

How @AlexandervonHumboldt2 explained to you.

OpenStudy (anonymous):

Okay thanks guys (:

OpenStudy (theedwardsfamily):

np! ^_^